Let P2 be the vector space of all polynomials of degree 2 or less, and let H be the subspace spanned by 8x−5x2+3, 2x-2x2+1 and 3x2-1.

a) The dimensions of the subspace H is ___________?

b) Is {8x-5x2+3, 2x-2x2+1, 3x2-1} a basis for P2? ________(be sure to explain and justify answer)

c) A basis for the subspace H is {_________}? enter a polynomial or comma separated list of polynomials
Answer #1

Now we see that. 8u 5m + 3 = 4(2x – 2x2+1) ++(342-1) 311 H = spom 80-5073, 20-2n#, 3mm (2n- 2 +) (3x+9 = spen . The dimension